George’s age increased by 15 is 7 less than 3 times his age. How old is George?
Answer by macston(5194)

G=George's age
G+15=3G-7 Add 7 to each side.
G+22=3G Subtract G from each side.
22=2G Divide each side by 2.
11=G ANSWER: George is 11.
CHECK
G+15=3G-7
11+15=3(11)-7
26=33-7
26=26